# Characterizing (multi-)semi-directed phylogenetic networks

### v1

@@ 4 Rootings of multi-semi-directed networks
The following theorem and its corollary show that the number of roots of a rooting of a multi-semi-directed network $N=(V,E,A)$ can be calculated directly from the reticulation number $|A|-|R|$ of $N$, with $R$ the set of reticulations.
@@
$|A|-|R|=|E|+|A|-|V|+k.$
Since $|A|-|R|$ is the reticulation number of $N$, the theorem follows. ∎
@@
For a rooting $D$ of multi-semi-directed network $N=(V,E,A)$, define the root locations as the triple $(V^{\prime},E^{\prime},A^{\prime})$ with $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$ and $A^{\prime}\subseteq A$ such that the roots of $D$ are precisely the vertices in $V^{\prime}$ together with vertices subdividing each edge in $E^{\prime}$ and arc in $A^{\prime}$. In the following theorem, we characterize the valid root locations of a multi-semi-directed network. As noted in the introduction, a similar characterization was given in (maxfield2025dissimilarity, , Proposition 8), although it was under a slightly different framework (e.g. assuming $A^{\prime}=E^{\prime}=\emptyset$).
@@
Let $N=(V,E,A)$ be a multi-semi-directed network and $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$ and $A^{\prime}\subseteq A$. Then there exists a rooting $D$ of $N$ with root locations $(V^{\prime},E^{\prime},A^{\prime})$ if and only if
@@
We first prove the “only if” direction. Since $N$ is multi-semi-directed, there exists a rooting $D$ of $N$. Let $(V^{\prime\prime},E^{\prime\prime},A^{\prime\prime})$ be its root locations.
@@
We now prove the “if” direction. Since $N$ is multi-semi-directed, there exists a rooting $D$ of $N$. Let $(V^{\prime\prime},E^{\prime\prime},A^{\prime\prime})$ be its root locations. For an edge $\{u,v\}$ of $N$, clearly, there exists a rooting of $N$ with root $u$ if and only if there exists a rooting with root $w$ subdividing $\{u,v\}$. Similarly, for an arc $(u,v)$ of $N$, clearly, there exists a rooting of $N$ with root $u$ if and only if there exists a rooting with root $w$ subdividing $(u,v)$. Hence, we may assume $E^{\prime}=A^{\prime}=E^{\prime\prime}=A^{\prime\prime}=\emptyset$. Let $R$ be the set of reticulations of $N$. Since each element of $V^{\prime}$ is in a different source component, there is no edge-path between any two vertices in $R\cup V^{\prime}$. Then, by Lemma 3, it follows that there exists a rooting of $N$ with root locations $(V^{\prime},E^{\prime},A^{\prime})$. ∎
Theorem 4 directly leads to an efficient algorithm for deciding if an arbitrary mixed graph $G$ is a (multi-)semi-directed network (note that in (maxfield2025dissimilarity, , Remark 1) a similar algorithm is sketched to efficiently find a rooting of a multi-semi-directed network). First, find all the source components of $G$. Second, in each source component, pick an arbitrary vertex as root. Third, do a breath-first search from each root and orient all edges away from the root. Finally, check if the resulting mixed graph $G^{\prime}$ is a multi-rooted directed network and whether its semi-deorientation is $G$. Note that by additionally checking whether or not $G^{\prime}$ has a single root, it can also be determined if $G$ is semi-directed.
@@ 5 Omnians
Figure 6: The three possible shapes of an omnian in a multi-semi-directed network: (i) one incident edge and at least two outgoing arcs, (ii) at least two incoming arcs and one outgoing arc and (iii) at least two outgoing arcs (and in all cases no other incident edge/arcs). Note that Case (iii) is possible in multi-semi-directed networks but not in semi-directed networks.
@@
We next turn our attention to tree-based multi-semi-directed networks. A spanning tree of a mixed graph $G$ is a subtree of $G$ that contains all vertices of $G$ and is a tree. A multi-rooted network $N$ is tree-based if it has a rooted spanning tree that has the same leaf set as $N$ (see semple2016phylogenetic ). A multi-semi-directed network $N$ is weakly tree-based if $N$ has a rooting that is tree-based and $N$ is strongly tree-based if every rooting of $N$ is tree-based, see Figure 8 for an example. Clearly, all multi-semi-directed networks that are weakly tree-based are semi-directed networks. Hence, we will focus on semi-directed networks for the remainder of this section.
@@ 6 Cherry Picking
To introduce cherry picking for multi-rooted networks, consider a multi-rooted network $D$ on $X$. A reticulated cherry of $D$ is an ordered pair of leaves $(x,y)$ such that $y$ is a reticulation leaf and there is a length-$3$ path between $x$ and $y$. Reducing a reticulated cherry $(x,y)$ of $D$ means deleting the arc from the parent of $x$ to the parent of $y$ and suppressing any resulting non-root degree-2 vertices. Let $D(x,y)$ be the result of reducing a cherry or reticulate cherry $(x,y)$ in $D$. Then the leaf set of $D(x,y)$ is either the same as the leaf set of $D$ or, if $(x,y)$ is a cherry of $D$, then the leaf set of $D(x,y)$ is $X\setminus\{x\}$. We say that $D$ is orchard if it can be reduced to a disjoint union of arcs using a sequence of cherry reductions and reticulated cherry reductions.
@@
Suppose that $N$ is a multi-semi-directed network. Consider an arbitrary rooting $D$ of $N$. If a root $\rho$ of $D$ is not a vertex of $N$, then it subdivides an edge or an arc of $N$. If $\rho$ subdivides $e$, we can make any vertex incident with $e$ a root instead of $\rho$. If $\rho$ subdivides an arc $a$ whose tail is not in $X$, then we can make the tail of $a$ a root instead of $\rho$. ∎
@@
    $N$ has a rooting that has a binary resolution that admits an HGT-consistent labelling;
@@
Figure 10: Example of a weakly orchard multi-semi-directed network on $\{a\ldots,e\}$ and a sequence of cherry reductions and reticulated cherry reductions. In each case, the (reticulated) cherry that is reduced is indicated below the arrow that indicates the reduction.
Figure 11: An alternative sequence of cherry reductions and reticulated cherry reductions for the weakly orchard multi-semi-directed network from Figure 10. In each case, the (reticulated) cherry that is reduced is indicated below the arrow that indicates the reduction.
@@
The other direction of the proof is by induction on $k=|X|+|R|-2$, with $R$ the set of reticulations of $N$. For $k=0$ the statement is trivially true. Assume $k\geq 1$ and that for each strong cherry picking sequence $s=(s_{1},\ldots,s_{k})$ of $N_{0}=N$ and for each $i\in\{1,\ldots,k\}$, it holds that if $s_{i}$ is an scr-cherry of $N_{i-1}=N\circ(s_{1},\ldots,s_{i-1})$ then $N_{i-1}$ has at least two scr-cherries.
@@
The example in Figure 12 shows why in the characterization in Theorem 8 “for each” cannot be replaced by “there exists”. Even though there exists a cherry picking sequence of the required type in the depicted semi-directed network $N$, $N$ is not strongly orchard. Indeed, $N$ also has a strong cherry picking sequence that is not of the required type.
Figure 12: A semi-directed network $N$ on $\{a,b,c,d,e\}$ that is weakly orchard but not strongly orchard, along with a rooting $D$ that is not orchard. This is easy to see since $D$ has no cherry, $(a,b)$ is the only reticulated cherry of $D$, and the depicted network $D(a,b)$ has no cherries or reticulated cherries. The bottom row of networks shows the first $4$ networks obtain when reducing $N$ by the strong cherry picking sequence $((d,e),(d,e),(c,e),(b,e),(a,b),(b,d))$ of $N$.
@@ 7 Path Partitions
We now turn to our final concept for characterizing classes of (multi-)semi-directed networks networks called path partitions. Basically speaking, for a multi-rooted network, this is a partition of the vertex set of the network whose parts induce a collection of directed paths each of which must end in a leaf. Multi-rooted networks that enjoy this property are called forest-based networks and can be used to model introgression scholz2019osf ; they are considered in more depth in lafond2025path ; huber2022forest . In this section, we extend the theory of path partitions from multi-rooted networks to (multi-)semi-directed networks and shall see that having a path partition characterizes weakly tree-based semi-directed networks (Corollary 4).
[…]
@@
    each semi-directed path in $\mathcal{P}$ is either a trivial path $(x)$ with $x\in X$ or a semi-directed path from a vertex in $V(N)-X$ to some element in $X$;
@@ Acknowledgements
This paper is based on research that was partly carried out while the authors were in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, RI, during the semester program on “Theory, Methods, and Applications of Quantitative Phylogenomics”. We thank the organizers for organizing this semester program. We thank Cécile Ané for posing an open question during this semester program regarding weakly and strongly tree-child semi-directed networks, which motivated part of this research.

### v2

@@ 4 Rootings of multi-semi-directed networks
The following theorem and its corollary show that the number of roots of a rooting of a multi-semi-directed network $N=(V,E,A)$ can be calculated directly from the reticulation number |A||R||A|-|R| of $N$, with $R$ the set of reticulations.
@@
|A||R|=|E|+|A||V|+k.|A|-|R|=|E|+|A|-|V|+k.
Since |A||R||A|-|R| is the reticulation number of $N$, the theorem follows. ∎
@@
For a rooting $D$ of multi-semi-directed network $N=(V,E,A)$, define the root configuration as the triple $(V^{\prime},E^{\prime},A^{\prime})$ with $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$ and $A^{\prime}\subseteq A$ such that the roots of $D$ are precisely the vertices in $V^{\prime}$ together with vertices subdividing each edge in $E^{\prime}$ and arc in $A^{\prime}$. In the following theorem, we characterize the valid root configurations of a multi-semi-directed network. As noted in the introduction, a similar characterization was given in (maxfield2025dissimilarity, , Proposition 8), although it was under a slightly different framework (e.g. assuming $A^{\prime}=E^{\prime}=\emptyset$).
@@
Let $N=(V,E,A)$ be a multi-semi-directed network and $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$ and $A^{\prime}\subseteq A$. Then there exists a rooting $D$ of $N$ with root configuration $(V^{\prime},E^{\prime},A^{\prime})$ if and only if
@@
We first prove the “only if” direction. Since $N$ is multi-semi-directed, there exists a rooting $D$ of $N$. Let $(V^{\prime\prime},E^{\prime\prime},A^{\prime\prime})$ be its root configuration.
@@
We now prove the “if” direction. Since $N$ is multi-semi-directed, there exists a rooting $D$ of $N$. Let $(V^{\prime\prime},E^{\prime\prime},A^{\prime\prime})$ be its root configuration. For an edge $\{u,v\}$ of $N$, clearly, there exists a rooting of $N$ with root $u$ if and only if there exists a rooting with root $w$ subdividing $\{u,v\}$. Similarly, for an arc $(u,v)$ of $N$, clearly, there exists a rooting of $N$ with root $u$ if and only if there exists a rooting with root $w$ subdividing $(u,v)$. Hence, we may assume $E^{\prime}=A^{\prime}=E^{\prime\prime}=A^{\prime\prime}=\emptyset$. Let $R$ be the set of reticulations of $N$. Since each element of $V^{\prime}$ is in a different source component, there is no edge-path between any two vertices in $R\cup V^{\prime}$. Then, by Lemma 3, it follows that there exists a rooting of $N$ with root configuration $(V^{\prime},E^{\prime},A^{\prime})$. ∎
Theorem 4 directly leads to an efficient algorithm taking $O(|V|+|E|+|A|)$ time for deciding if an arbitrary mixed graph $G=(V,E,A)$ is a (multi-)semi-directed network (note that in (maxfield2025dissimilarity, , Remark 1) a similar algorithm is sketched to efficiently find a rooting of a multi-semi-directed network). First, find all the source components of $G$ by traversing the graph in $O(|V|+|E|+|A|)$ time. Second, in each source component, pick an arbitrary vertex as root. Third, do a breath-first search from each root and orient all edges away from the root, again taking $O(|V|+|E|+|A|)$ time. Finally, with yet another traversal, check if the resulting mixed graph $G^{\prime}$ is a multi-rooted directed network and whether its semi-deorientation is $G$. Note that by additionally checking whether or not $G^{\prime}$ has a single root, it can also be determined if $G$ is semi-directed.
@@ 5 Omnians
Figure 6: The three possible configurations surrounding an omnian in a multi-semi-directed network: (i) one incident edge and at least two outgoing arcs, (ii) at least two incoming arcs and one outgoing arc and (iii) at least two outgoing arcs (and in all cases no other incident edge/arcs). Note that Case (iii) is possible in multi-semi-directed networks but not in semi-directed networks.
@@
We next turn our attention to tree-based multi-semi-directed networks. A spanning tree of a mixed graph $G$ is a subtree of $G$ that contains all vertices of $G$ and is a tree. A multi-rooted network $N$ is tree-based if it has a rooted spanning tree that has the same leaf set as $N$. Clearly, a multi-rooted network that is tree-based must be a rooted network, since the corresponding spanning tree must have a single root. Note that binary rooted tree-based networks were introduced in francis2015which and extended to non-binary rooted networks in jetten2016nonbinary , who also defined the stricter notion of strictly tree-basedness, which we do not consider here. A multi-semi-directed network $N$ is weakly tree-based if $N$ has a rooting that is tree-based and $N$ is strongly tree-based if every rooting of $N$ is tree-based, see Figure 8 for an example. Since every multi-rooted tree-based network is a rooted network, all multi-semi-directed networks that are weakly or strongly tree-based are semi-directed networks. Hence, we will focus on semi-directed networks for the remainder of this section.
@@ 6 Cherry Picking
To introduce cherry picking for multi-rooted networks, consider a multi-rooted network $D$ on $X$. A reticulated cherry [^5] Note that this definition is only for multi-rooted networks. Reticulated cherries in multi-semi-directed networks are defined in a slightly different way below. of $D$ is an ordered pair of leaves $(x,y)$ such that $y$ is a reticulation leaf and there is a length-$3$ path between $x$ and $y$. Reducing a reticulated cherry $(x,y)$ of $D$ means deleting the arc from the parent of $x$ to the parent of $y$ and suppressing any resulting non-root degree-2 vertices. Let $D(x,y)$ be the result of reducing a cherry or reticulate cherry $(x,y)$ in $D$. Then the leaf set of $D(x,y)$ is either the same as the leaf set of $D$, or, if $(x,y)$ is a cherry of $D$, then the leaf set of $D(x,y)$ is $X\setminus\{x\}$. We say that $D$ is orchard if it can be reduced to a disjoint union of arcs using a sequence of cherry reductions and reticulated cherry reductions.
@@
Suppose that $N$ is a multi-semi-directed network. Consider an arbitrary rooting $D$ of $N$. If a root $\rho$ of $D$ is not a vertex of $N$, then it subdivides an edge $e$ or an arc $a$ of $N$. If $\rho$ subdivides $e$, we can make any vertex incident with $e$ a root instead of $\rho$. If $\rho$ subdivides an arc $a$ whose tail is not in $X$, then we can make the tail of $a$ a root instead of $\rho$. ∎
@@
    NN has a rooting that has a binary resolution that admits an HGT-consistent labelling;
@@
Figure 10: Example of a weakly orchard multi-semi-directed network on $X=\{a\ldots,e\}$ and a sequence of cherry reductions and reticulated cherry reductions. In each case, the (reticulated) cherry that is reduced is indicated below the arrow that indicates the reduction.
Figure 11: An alternative sequence of cherry reductions and reticulated cherry reductions for the weakly orchard multi-semi-directed network on $X=\{a\ldots,e\}$ from Figure 10. In each case, the (reticulated) cherry that is reduced is indicated below the arrow that indicates the reduction. Note that in the bottom-left multi-semi-directed network, $a$ has become a root instead of a leaf and that $(a,d)$ is a reticulated cherry in that network since $a\in X$.
@@
Remark.
It follows from Theorem 7, Lemmas 4 and 5, and Figure 12 that there is a key distinction between the weakly and strongly orchard properties concerning the reduction of (reticulated) cherries in arbitrary order. Specifically, the property of a multi-semi-directed network being weakly orchard is preserved under arbitrarily reducing both cherries and reticulated cherries. In contrast, a semi-directed network being strongly orchard is preserved only under arbitrarily reducing cherries and reticulated cherries that are not scr-cherries.
@@
The other direction of the proof is by induction on k=|X|+|R|2k=|X|+|R|-2, with $R$ the set of reticulations of $N$. For $k=0$ the statement is trivially true. Assume $k\geq 1$ and that for each strong cherry picking sequence $s=(s_{1},\ldots,s_{k})$ of $N_{0}=N$ and for each $i\in\{1,\ldots,k\}$, it holds that if $s_{i}$ is an scr-cherry of $N_{i-1}=N\circ(s_{1},\ldots,s_{i-1})$ then $N_{i-1}$ has at least two scr-cherries.
@@
The example in Figure 12 shows why in the characterization in Theorem 8 “for each” cannot be replaced by “there exists”. Even though there exists a cherry picking sequence of the required type in the depicted semi-directed network $N$, NN is not strongly orchard. Indeed, $N$ also has a strong cherry picking sequence that is not of the required type.
Figure 12: A semi-directed network $N$ on $\{a,b,c,d,e\}$ that is weakly orchard but not strongly orchard, along with a rooting $D$ that is not orchard. This is easy to see since $D$ has no cherry, $(a,b)$ is the only reticulated cherry of $D$, and the depicted network $D(a,b)$ has no cherries or reticulated cherries. The bottom row of networks shows the first 44 networks obtain when reducing $N$ by the strong cherry picking sequence $((d,e),(d,e),(c,e),(b,e),(a,b),(b,d))$ of $N$.
@@ 7 Path Partitions
We now turn to our final concept for characterizing classes of (multi-)semi-directed networks networks called path partitions, see also francis2018new ; lafond2025path ; huber2022forest . Basically speaking, for a multi-rooted network, this is a partition of the vertex set of the network whose parts induce a collection of directed paths each of which must end in a leaf. Multi-rooted networks that enjoy this property are called forest-based networks and can be used to model introgression scholz2019osf ; they are considered in more depth in lafond2025path ; huber2022forest . In this section, we extend the theory of path partitions from multi-rooted networks to (multi-)semi-directed networks and shall see that having a path partition characterizes weakly tree-based semi-directed networks (Corollary 4).
@@
    each semi-directed path in $\mathcal{P}$ is either a trivial path $(x)$ with $x\in X$ or a semi-directed path from a vertex in V(N)XV(N)-X to some element in $X$;
@@ Acknowledgements
This paper is based on research that was partly carried out while the authors were in residence at the Institute for Computational and Experimental Research in Mathematics (ICERM) in Providence, RI, during the semester program on “Theory, Methods, and Applications of Quantitative Phylogenomics”. We thank the organizers for organizing this semester program. We thank Cécile Ané for posing an open question during this semester program regarding weakly and strongly tree-child semi-directed networks, which motivated part of this research. We thank the reviewers for their valuable comments and suggestions.